# High-order WENO finite-difference methods for hyperbolic nonconservative systems of Partial Differential Equations

### v1

@@ Abstract
This work aims to extend the well-known high-order WENO finite-difference methods for systems of conservation laws to nonconservative hyperbolic systems. The main difficulty of these systems both from the theoretical and the numerical points of view comes from the fact that the definition of weak solution is not unique: according to the theory developed by Dal Maso, LeFloch, and Murat in 1995, it depends on the choice of a family of paths. A general strategy is proposed here in which WENO operators are not only used to reconstruct fluxes but also the nonconservative products of the system. Moreover, if a Roe linearization is available, the nonconservative products can be computed through matrix-vector operations instead of path-integrals. The methods are extended to problems with source terms and two different strategies are introduced to obtain well-balanced schemes. These numerical schemes will be then applied to the two-layer shallow water equations in one- and two- dimensions to obtain high-order methods that preserve water-at-rest steady states.
@@ 1 Introduction
will be also considered.
@@
where $\sigma$ is the propagation speed; $U^{\pm}$ the lateral limits of the solution; and $\vec{n}=(n_{1},n_{2})$ a unit vector normal to the jump.
Based on this theory, a framework to design finite-volume methods for nonconservative sysmes was introduced in [37] based on the concept of path-conservative methods. These methods have been extensively applied to solve nonconservative systems: see, for instance, [8, 11, 13, 16, 17, 19].
In this paper, we focus on WENO finite-difference schemes. In the past decades, these methods have been widely applied to systems of conservation and balance laws: see for instance [5, 7, 28, 29, 43, 44, 48]. In [22] a fifth-order A-WENO finite-difference scheme for 1D and 2D systems of nonconservative hyperbolic systems was introduced. This scheme is based on the path-conservative central-upwind method and the global flux approach: the integral of the nonconservative term is considered as a new flux function of the system.
The goal of this paper is to introduce a new approach to design high-order finite-difference methods for nonconservative systems. The key idea is to apply a standard high-order WENO reconstruction operator to the nonconservative products computed using the selected family of paths: more precisely, in the 1D case, the selected WENO operator will be applied to reconstruct quantities of the form
@@
where $\mathcal{S}_{i}$ represents the stencil of the $i$th point and $A^{\pm}(U)$ represents a splitting of the matrix system $A(U)$. Two different splittings will be considered here: Lax-Friedrichs and upwind splittings. It will be seen that, if a path-consistent Roe linearization is available (see [15]), the quantities to be reconstructed can be computed by using matrix-vector products instead of path-integrals. Compared to the approach in [22], the main advantages of this new one are the following:
  • the accuracy in space of the methods only depends on the order of the standard WENO reconstruction selected: no order-dependent corrections of the fluxes are necessary;
  • no integrals involving quadrature points in the cells have to be computed which avoids having to use an additional reconstruction operator with uniform accuracy in the entire cells.
The numerical methods will be applied to the 1D and 2D hyperbolic two-layer shallow-water equations that govern the flow of two superposed layers of immiscible homogeneous fluids. This system is used in different ocean and coastal engineering simulations and there is a vast literature focusing on its numerical analysis: see for instance [10, 12, 21, 22, 23, 33, 36].
A relevant property to be satisfied by the numerical methods solving systems of the form (1) or (3) is the preservation of some or all the steady-state solutions of the system, i.e. the well-balanced property. For instance, in the context of the one or the two-layer shallow-water equations, a minimal requirement to the numerical methods is to exactly preserve the steady states corresponding to water-at-rest, i.e. to satisfy the C-property according to [4]. Different techniques have been proposed to design well-balanced schemes including hydrostatic reconstruction related [2, 18], relaxation methods [1], consistent discretization of the flux and source term [32, 46, 47], etc. See also [6, 26, 34].
In the context of finite-volume methods, a general strategy to design well-balanced methods has been described in [20]. In this strategy, a stationary solution whose average is the numerical approximation at every cell has to be computed at every time step. Then, a standard reconstruction operator is applied to the differences of the cell values at the stencil and the cell averages of the local stationary solution. In [38] this strategy has been extended to WENO finite-difference methods for systems of balance laws. Two different strategies will be followed here to obtain well-balanced numerical methods: one of them is the extension to nonconservative systems of the strategy introduced in this last reference, while the other consists in combining the upwind splitting with an adequate choice of family of paths.
@@
The rest of the paper is organized as follows. In Section 2, firstly, the new WENO path-conservative schemes for 1D homogeneous (i.e. without source terms) nonconservative systems are introduced and their high-order accuracy property is proved. In Section 3, source terms are included and two strategies to obtain well-balanced methods are described. In Section 4, the proposed schemes are extended to 2D nonconservative systems. In Section 5, we apply the proposed scheme to 1D and 2D two-layer shallow water equations: the numerical results are presented in Section 6. Finally, some conclusions are drawn in Section 7.
@@ 2 Path-Conservative WENO finite-difference reconstruction methods
2.1 WENO methods
@@
It will be assumed here that the system is strictily hyperbolic, i.e. for every $U$, the matrix $A(U)$ has N different real eigenvalues
@@
where $F_{i+1/2}$ is a high-order reconstruction of the flux function. Uniform meshes of constant step $\Delta x$ will be considered and the following notation will be also used for the intercells
$x_{i+1/2}=x_{i}+\frac{\Delta x}{2}.$
@@
These are the so-called left- and right-biased reconstructions, related throught the equality:
@@
$U_{i}^{\prime}+\frac{1}{\Delta x}\left(\widehat{D}^{-}_{i+1/2}+\widehat{D}^{+}_{i-1/2}\right)=0,$ (14)
@@
In order to extend the method to (5) first a splitting of $A(U)$
$A(U)=A^{+}(U)+A^{-}(U),$
has to be chosen in such a way that the eigenvalues of $A^{+}(U)$ (resp. $A^{-}(U)$) are positive (resp. negative). For instance, the LF splitting can be extended in a natural manner:
$A^{\pm}(U)=\frac{1}{2}\left(A(U)\pm\alpha I\right),$
with $\alpha$ is computed as in the conservative case.
Finally, the last ingredient required to extend the numerical methods to nonconservative systems is a family of paths $\Psi:[0,1]\times\Omega\times\Omega\to\Omega$. Using $\Psi$ we have:
$D^{\pm}_{j,k}=F^{\pm}(U_{k})-F^{\pm}(U_{j})=\int_{0}^{1}DF^{\pm}(\Psi(s;U_{j},U_{k}))\partial_{s}\Psi(s;U_{j},U_{k})\,ds.$
The natural extension of the numerical method (7) to the system (5) is then given by (14) with
@@
$D^{\pm}_{j,k}=\int_{0}^{1}A^{\pm}(\Psi(s;U_{j},U_{k}))\partial_{s}\Psi(s;U_{j},U_{k})\,ds.$ (17)
Observe that, while in the case of a system of conservation laws, the resulting numerical method is independent of the chosen family of paths (since it is equivalent to (7)), for nonconservative systems the numerical method depends on the chosen family of paths. Please note that, while for systems of conservation laws only two reconstructions per intercell are required, here 4 reconstructions are needed in $x_{i+1/2}$: two reconsttructions to compute $\widehat{D}^{+}_{i+1/2}$ and two others to compute $\widehat{D}^{-}_{i+1/2}$
Let us suppose that a Roe linearization is available for the chosen family of paths $\Psi$, i.e. a matrix-valued function $A_{\Psi}:\Omega\times\Omega\mapsto\mathbb{R}^{N}\times\mathbb{R}^{N}$ that satisfies the following properties:
@@
Then, the fluctuations $D_{j,k}$ can be computed as follows:
@@
where $A^{\pm}_{\Psi}\left(U,V\right)$ represents the application to $A_{\Psi}\left(U,V\right)$ of the splitting chosen for $A(U)$: for instance, if the LF splitting has been chosen, then
$A^{\pm}_{\Psi}\left(U,V\right)=\frac{1}{2}\left(A_{\Psi}\left(U,V\right)\pm\alpha I\right).$
It is well-known fact that high-order finite-difference methods are less oscillatory if the reconstructions are performed in characteristic fields. To do that, let us suppose that a Roe linearization $A_{\Psi}$ is available. Given two indices $j,k$ the following notation will be used
$A_{j,k}=A_{\Psi}(U_{j},U_{k})$
to represent the Roe matrix. This matrix can be decomposed in the form:
$A_{j,k}={L}_{j,k}\Lambda_{j,k}{R}_{j,k},$
where $\Lambda_{j,k}$ is the $N$-dimensional diagonal matrix whose coefficients are the eigenvalues of $A_{j,k}$:
$\lambda_{j,k;1},\dots,\lambda_{j,k;N};$
$R_{j,k}$ is a matrix whose $l$th column $\vec{r}_{j,k;l}$ is an eigenvector associated to $\lambda_{j,k;l}$; and ${L}_{j,k}={R_{j,k}^{-1}}$ is a matrix whose arrows are left-eigenvalues. To compute the reconstruction in characteristic variables, the fluctuations involved in the computation of $\widehat{D}^{\pm}_{i+1/2}$ are first multiplied by the matrix $L_{i,i+1}$ of left eigenvectors of $A_{i,i+1}$ and the reconstructed fluxes are then multiplied by the matrix $R_{i,i+1}$ of right eigenvectors, i.e.
@@
An alternative implementation of WENO methods is given by the upwind approach, in which the splitting of the Roe matrix is given by:
$A^{\pm}_{\Psi}\left(U,V\right)=P^{\pm}_{\Psi}\left(U,V\right)A_{\Psi}\left(U,V\right).$
Here, $P^{\pm}$ represent the projection matrices onto the positive/negative eigenvectors of the Roe matrix, i.e.
$P^{\pm}_{\Psi}\left(U,V\right)=R_{\Psi}\left(U,V\right)M^{\pm}_{\Psi}\left(U,V\right)R^{-1}_{\Psi}\left(U,V\right),$
where $M^{\pm}_{\Psi}\left(U,V\right)$ is the diagonal matrix whose coefficients are
$\frac{1}{2}\left(1\pm\operatorname{sign}\left(\lambda_{j}(U,V)\right)\right),\quad j=1,\ldots,N,$
and $R_{\Psi}\left(U,V\right)$ is a matrix whose $j$th columns is an eigenvector associated to $\lambda_{j}(U,V)$. This strategy is implemented as follows: given two indices $j,k$, let $\{\alpha_{j,k;l}\}_{l=1}^{N}$ be the coordinates of $U_{k}-U_{j}$ in the basis of eigenvectors of the Roe matrix $A_{j,k}$, i.e.
$U_{k}-U_{j}=\sum_{l=1}^{N}\alpha_{j,k;l}\vec{r}_{j,k;l}.$
Then
$D^{\pm}_{j,k}=A^{\pm}_{j,k}(U_{k}-U_{j})=\sum_{l=1}^{N}\alpha_{j,k;l}\lambda^{\pm}_{j,k;l}\vec{r}_{j,k;l},$
$\lambda^{+}=\frac{\lambda+|\lambda|}{2},\quad\lambda^{-}=\frac{\lambda-|\lambda|}{2}.$
Observe that, if the LF splitting is chosen and reconstruction in characteristic variables is performed, the right and left eigenvectors of every Roe matrix have to be computed. On the other hand, if the upwind approach is followed, the eigenvectors and eigenvalues of the Roe matrices $A_{k,j}$ are required. Moreover, the linear system
$R_{j,k}\vec{\alpha}_{j,k}=U_{k}-U_{j},$
has to be solved to compute the coordinates $\alpha_{j,k;l}$. Therefore, it is more computationally expensive.
Remark.
@@
$U_{i}^{\prime}=-\frac{1}{\Delta x}\left(F_{i+1/2}-F_{i-1/2}+\widehat{B}^{-}_{i+1/2}+\widehat{B}^{+}_{i-1/2}\right),$ (22)
where $F_{i+1/2}$ and $\widehat{B}_{i+1/2}$ are, respecitvely, standard WENO reconstructions of the flux function and the nonconservative terms, in which the nonlinear coefficients are the same.

### v2

@@ Abstract
This work aims to extend the well-known high-order WENO finite-difference methods for systems of conservation laws to nonconservative hyperbolic systems. The main difficulty of these systems both from the theoretical and the numerical points of view comes from the fact that the definition of weak solution is not unique: according to the theory developed by Dal Maso, LeFloch, and Murat in 1995, it depends on the choice of a family of paths. A new strategy is introduced here that allows non-conservative products to be written as the derivative of a generalized flux function that is defined locally on the basis of the selected family of paths. WENO reconstructions are then applied to this generalized flux. Moreover, if a Roe linearization is available, the generalized flux function can be evaluated through matrix-vector operations instead of path-integrals. Two different known techniques are used to extend the methods to problems with source terms and the well-balanced properties of the resulting schemes are studied. These numerical schemes are applied to a coupled Burgers’ system and to the two-layer shallow water equations in one- and two- dimensions to obtain high-order methods that preserve water-at-rest steady states.
@@ 1 Introduction
will also be considered.
@@
The numerical methods will be applied to the 1D and 2D hyperbolic two-layer shallow-water equations that govern the flow of two superposed layers of immiscible homogeneous fluids. This system is used in different ocean and coastal engineering simulations and there is a vast literature focusing on its numerical analysis: see for instance [14, 17, 25, 26, 27, 38, 41]. In order to illustrate the general procedure, we will also consider the 1D coupled Burgers’ system introduced in [19]:
$\left\{\begin{array}[]{l}u_{t}+uu_{x}+uv_{x}=0,\\ v_{t}+vu_{x}+vv_{x}=0,\end{array}\right.$ (4)
that can be written in the form (2) with
$A(U)=\left[\begin{array}[]{cc}u&u\\ v&v\end{array}\right],\quad U\in\Omega,$
with
$\Omega=\left\{\left[\begin{array}[]{c}u\\ v\end{array}\right]\in\mathbb{R}^{2}\text{ s.t. }u+v>0\right\}.$
The eigenvalues of $A(U)$ are
$\lambda_{1}(U)=0,\quad\lambda_{2}(U)=u+v,$
and as so that the system is strictly hyperbolic in $\Omega$.
@@
where $\sigma$ is the propagation speed; $U^{\pm}$ the lateral limits of the solution; and $\vec{n}=(n_{1},n_{2})$ a unit vector normal to the jump. For instance, in the particular case of system (4), it can be checked that the choice of the family of straight segments
$\Psi_{1}(s;U^{-},U^{+})=U^{-}+s(U^{+}-U^{-}),\quad s\in[0,1]$ (5)
leads to the jump conditions
$\left\{\begin{array}[]{l}\displaystyle\bar{u}(u^{+}-u^{-})+\bar{u}(v^{+}-v^{-})=\sigma(u^{+}-u^{-}),\\ \displaystyle\bar{v}(u^{+}-u^{-})+\bar{v}(v^{+}-v^{-})=\sigma(v^{+}-v^{-}),\end{array}\right.$ (6)
with
$\bar{u}=\frac{u^{+}+u^{-}}{2},\quad\bar{v}=\frac{v^{+}+v^{-}}{2},$
while the choice
$\Psi_{2}(s;U^{-},U^{+})=\left[\begin{array}[]{c}u^{-}+s(4-3s)(u^{+}-u^{-})\\ v^{-}+s(v^{+}-v^{-})\\ \end{array}\right],\quad s\in[0,1],$ (7)
leads to
$\left\{\begin{array}[]{c}\displaystyle\bar{u}(u^{+}-u^{-})+u^{+}(v^{+}-v^{-})=\sigma(u^{+}-u^{-}),\\ \displaystyle v^{-}(u^{+}-u^{-})+\bar{v}(v^{+}-v^{-})=\sigma(v^{+}-v^{-}).\end{array}\right.$ (8)
Any choice of family paths leads to a consistent definition of weak solutions from the mathematical point of view. Therefore, given a particular application, the choice of the adequate family of paths has to be based on the consistency with the physics of the problem: for instance, the family of paths can be given by the viscous profiles corresponding to the neglected viscous terms (see [15] for a general discussion and [7] for the particular case of System (4)).
Based on this theory, a framework for designing finite-volume methods for nonconservative systems was introduced in [42] based on the concept of path-conservative methods. These methods have been extensively applied to solve nonconservative systems: see, for instance, [12, 16, 20, 21, 23].
In this paper, we focus on WENO finite-difference schemes. In the past decades, these methods have been widely applied to systems of conservation and balance laws: see for instance [8, 11, 32, 33, 49, 53]. In [26] a fifth-order A-WENO finite-difference scheme for 1D and 2D systems of nonconservative hyperbolic systems was introduced. This scheme is based on the path-conservative central-upwind method and the global flux approach in which the integral of the nonconservative term is considered as a new flux function of the system. A high-order conservative method is then applied to the formal system of conservation laws. The evaluation of the new flux function requires the computation of the cell integrals of the nonconservative products using a high-order quadrature formula. Therefore, in addition to the flux reconstruction of the flux, it also requires high-order reconstructions of $U$ at the quadrature points. If this approach is used, the method is reduced to the conservative one obviously when $A(U)$ is the Jacobian of a flux function $F$. Recently, in [3] a finite-difference WENO method has been introduced based on state reconstruction that does not require the computation of cell integrals of the conservative products. In contrast to global flux-based methods, they reduce to the conservative method only when $A(U)$ is linear.
The goal of this paper is to introduce a local flux approach to design high-order finite-difference methods for nonconservative systems. The key idea is to apply a standard high-order WENO reconstruction operator to the nonconservative products computed using the selected family of paths: more precisely, in the 1D case, the selected WENO operator will be applied to reconstruct quantities of the form
@@
where $\mathcal{S}_{i}$ represents the stencil of the $i$th point and $A^{\pm}(U)$ represents a splitting of the matrix system $A(U)$. It will be seen that, if a path-consistent Roe linearization is available (see [47, 51]), the quantities to be reconstructed can be computed by using matrix-vector products instead of path-integrals. The main advantages of this new one are the following:
  • the accuracy in space of the methods only depends on the order of the selected WENO reconstruction provided that the selected family of paths satisfies a symmetry property to be described;
  • no integrals involving quadrature points in the cells have to be computed which avoids having to use an additional reconstruction operator with uniform accuracy in the entire cells;
  • the methods reduce to the standard finite-difference conservative WENO schemes when $A(U)$ is the Jacobian matrix of a flux function $F(U)$.
Two different matrix splittings will be considered here based on the standard Lax-Friedrichs and Upwind approaches to illustrate the strategy. Nevertheless it can also be applied to more general splittings or even to WENO reconstructions that are not based on splitting technique, as in the case of A-WENO. The application of the local flux approach combined with A-WENO reconstructions will be discussed in a forthcoming work.
A relevant property to be satisfied by the numerical methods solving systems of the form (1) or (3) is the preservation of some or all the steady-state solutions of the system, i.e. the well-balanced property. For instance, in the context of the one or the two-layer shallow-water equations, a minimal requirement to the numerical methods is to exactly preserve the steady states corresponding to water-at-rest, i.e. to satisfy the C-property according to [6]. Different techniques have been proposed to design well-balanced schemes including hydrostatic reconstruction related [2, 22], relaxation methods [1], consistent discretization of the flux and source term [36, 52], etc. See also [9, 29, 39].
In the context of finite-volume methods, a general strategy to design well-balanced methods has been described in [24]. In this strategy, a stationary solution whose average is the numerical approximation at every cell has to be computed at every time step. Then, a standard reconstruction operator is applied to the differences in the cell values at the stencil and the cell averages of the local stationary solution. In [44] this strategy has been extended to WENO finite-difference methods for systems of balance laws. Two different strategies will be followed here to obtain well-balanced numerical methods: one of them is the extension to nonconservative systems of the strategy introduced in this last reference, while the other consists of combining the Upwind splitting scheme with an adequate choice of family of paths.
@@
The rest of the paper is organized as follows. In Section 2, firstly, the new WENO path-conservative schemes for 1D homogeneous (i.e. without source terms) nonconservative systems are introduced and their high-order accuracy property is proved. In this section the general problem is considered, so that the family of paths is, in principle, arbitrary. Nevertheless, it will be shown that a symmetry property has to be satisfied to ensure the high-order accuracy of the method. In Section 3, source terms are included and two strategies to obtain well-balanced methods are described. In Section 4, the proposed schemes are extended to 2D nonconservative systems. In Section 5, we apply the proposed scheme to 1D and 2D two-layer shallow water equations: the numerical results are presented in Section 6. Finally, some conclusions are drawn in Section 7.
@@ 2 Path-Conservative WENO finite-difference reconstruction methods
2.1 Conservative WENO finite difference schemes: a brief overview
@@
It will be assumed here that the system is strictly hyperbolic, i.e. for every $U$, the matrix $A(U)$ has N different real eigenvalues
@@
where $F_{i+1/2}$ is a high-order reconstruction of the flux function. The computational domain is $[a,b]$. Uniform meshes with a constant step size $\Delta x$ will be considered, with cell centers denoted as $x_{i}$. The following notation is used for the cell interface:
$x_{i+\frac{1}{2}}=x_{i}+\frac{\Delta x}{2}.$
@@
These are the so-called left- and right-biased reconstructions, related by the equality:
@@
Any version of WENO reconstructions can be selected. In particular, in the numerical tests shown in Section 6, WENOZ is used (see [11]): for the sake of completeness, the expression of the fifth-order WENOZ reconstruction is recalled in Appendix A.
2.2 Extension to nonconservative systems
@@
$\frac{dU_{i}}{dt}+\frac{1}{\Delta x}\left(\widehat{D}^{-}_{i+1/2}+\widehat{D}^{+}_{i-1/2}\right)=0,$ (19)
@@
Finally, the last ingredient required to extend the numerical methods to nonconservative systems is a family of paths. Let us consider, in principle, an arbitrary family $\Psi:[0,1]\times\Omega\times\Omega\to\Omega$. Using $\Psi$ we have:
$D^{\pm}_{j,k}=F^{\pm}(U_{k})-F^{\pm}(U_{j})=\int_{0}^{1}JF^{\pm}\left(\Psi(s;U_{j},U_{k})\right)\partial_{s}\Psi(s;U_{j},U_{k})\,ds,$
where $JF^{\pm}$ represent the Jacobian matrices of $F^{\pm}$. The natural extension of the numerical method (12) to the system (10) is then given by (19) with
@@
$D^{\pm}_{j,k}=\int_{0}^{1}A^{\pm}(\Psi(s;U_{j},U_{k}))\partial_{s}\Psi(s;U_{j},U_{k})\,ds,$ (22)
where $A^{\pm}(\Psi(s;U_{j},U_{k})$ is a matrix-splitting to be adequately chosen. For instance, for system (4), since the two eigenvalues are positive, a natural choice is given by
$A^{+}(\Psi(s;U_{j},U_{k})=A(\Psi(s;U_{j},U_{k}),\quad A^{-}(\Psi(s;U_{j},U_{k})=0.$ (23)
Please note that, while for systems of conservation laws only two reconstructions per intercell are required, here 4 reconstructions are needed in $x_{i+1/2}$: two reconstructions to compute $\widehat{D}^{+}_{i+1/2}$ and two others to compute $\widehat{D}^{-}_{i+1/2}$.
Observe that, while in the case of a system of conservation laws, the resulting numerical method is independent of the chosen family of paths (since it is equivalent to (12)), for nonconservative systems the numerical method depends on the chosen family of paths. For instance, in the particular case of System (4), if the matrix-splitting is given by (23), the choice of the family of paths (5), whose corresponding jump condition is (6), leads to
$D^{+}_{j,k}=\left[\begin{array}[]{c}\displaystyle\bar{u}_{j,k}(u_{k}-u_{j})+\bar{u}_{j,k}(v_{k}-v_{j})\\ \displaystyle\bar{v}_{j,k}(u_{k}-u_{j})+\bar{v}_{j,k}(v_{k}-v_{j})\end{array}\right],\quad D^{-}_{j,k}=0,$ (24)
with
$\bar{u}_{j,k}=\frac{u_{j}+u_{k}}{2},\quad\bar{v}_{j,k}=\frac{v_{j}+v_{k}}{2},$
while the choice (7), whose corresponding jump condition is (8), leads to
$D^{+}_{j,k}=\left[\begin{array}[]{c}\displaystyle\bar{u}_{j,k}(u_{k}-u_{j})+u_{k}(v_{k}-v_{j})\\ \displaystyle v_{j}(u_{k}-u_{j})+\bar{v}_{j,k}(v_{k}-v_{j})\end{array}\right],\quad D^{-}_{j,k}=0.$ (25)
Definitions (5) and (7) lead to different numerical results: both of them are convergent for smooth solutions but, as it will be seen in Section 2.3, only (24) leads to a high-order accurate method. On the other hand, the two methods are expected to give different results for discontinuous solutions, since they are formally consistent with the different jump conditions, (6) or (8), corresponding to the selected paths. According to [18] this formal consistency does not ensure that the limits of the numerical solutions satisfy the expected jump conditions. In fact, the methods introduced here can fail in capturing correctly the discontinuities, as any other standard finite-difference type method. Nevertheless, they can be combined with techniques like the ones recently developed in [45, 46] to improve their convergence to the sought weak solutions.
While for (4) the fluctuations (22) can be easily computed, this may be more difficult in other cases where numerical quadrature can be used to compute the integrals. Nevertheless, an alternative form of the method can be given if a Roe linearization is available in which the path-integrals are replaced by matrix-vector products. Remember that a Roe linearization (see [47, 51]) is a matrix-valued function $A_{\Psi}:\Omega\times\Omega\mapsto\mathbb{R}^{N}\times\mathbb{R}^{N}$ that satisfies the following properties:
@@
For instance, it can be easily checked that the matrices
$A_{\Psi_{1}}(U^{-},U^{+})=\left[\begin{array}[]{cc}\bar{u}&\bar{u}\\ \bar{v}&\bar{v}\end{array}\right],\quad A_{\Psi_{2}}(U^{-},U^{+})=\left[\begin{array}[]{cc}\bar{u}&u^{+}\\ v^{-}&\bar{v}\end{array}\right],$ (27)
are Roe linearizations for system (4) related to the family of paths $\Psi_{1}$ and $\Psi_{2}$ given by (5) and (7) respectively. If a Roe linearization is available (as it is the case for the two-layer shallow-water system if the family of straight segments is selected) the fluctuations $D_{j,k}^{\pm}$ can be computed as in [21]:
@@
where
$A^{\pm}_{\Psi}\left(U_{j},U_{k}\right)=\frac{1}{2}\left(A_{\Psi}\left(U_{j},U_{k}\right)\pm Q_{\Psi}\left(U_{j},U_{k}\right)\right)$ (29)
represents a splitting of the Roe linearization. Two different splittings will be considered here:
  • Upwind splitting:
    $Q_{\Psi}\left(U_{j},U_{k}\right)=\left|A_{j,k}\right|,$ (30)
    where
    $\left|A_{j,k}\right|={R}_{j,k}\left|\Lambda_{j,k}\right|{L}_{j,k}.$
    Here, $\left|\Lambda_{j,k}\right|$ is the $N$-dimensional diagonal matrix whose coefficients are the absolute values of the eigenvalues of $A_{j,k}$:
    $|\lambda_{j,k;1}|,\dots,|\lambda_{j,k;N}|;$
    $R_{j,k}$ is a matrix whose $l$th column $\vec{r}_{j,k;l}$ is an eigenvector associated to $\lambda_{j,k;l}$; and ${L}_{j,k}={R_{j,k}^{-1}}$ is a matrix whose arrows are left-eigenvalues. A standard entropy-fix can be used to avoid the appearance of non-entropy discontinuities, like considering a regularization $|\cdot|_{\epsilon}$ of the absolute value function like in [31, 50].
  • Lax-Friedrichs(LF) splitting:
    $Q_{\Psi}\left(U_{j},U_{k}\right)=\alpha I,$ (31)
    where $I$ is the identity matrix and $\alpha$ is the global maximum of the absolute value of the eigenvalues, $\alpha\geq\lvert\lambda_{j,k;l}\rvert,\;l=1,\dots,N$.
The matrices involved in the Upwind splitting can be equivalently written as follows
$A_{\Psi}^{\pm}\left(U,V\right)=P_{\Psi}^{\pm}\left(U,V\right)A_{\Psi}\left(U,V\right),$ (32)
where
$P^{\pm}_{\Psi}\left(U,V\right)=R_{\Psi}\left(U,V\right)M^{\pm}_{\Psi}\left(U,V\right)R^{-1}_{\Psi}\left(U,V\right).$ (33)
Here $M^{\pm}_{\Psi}\left(U,V\right)$ represents the diagonal matrix whose coefficients are
$\frac{1}{2}\left(1\pm\operatorname{sign}\left(\lambda_{l}(U,V)\right)\right),\quad l=1,\ldots,N,$
and $R_{\Psi}\left(U,V\right)$ is a matrix whose $l$th columns is an eigenvector $\vec{r}_{l}\left(U,V\right)$ associated to $\lambda_{l}(U,V)$. The fluctuations corresponding to this splitting can be then computed as follows: given two indices $j,k$, first the coordinates $\{\alpha_{j,k;l}\}_{l=1}^{N}$ of $U_{k}-U_{j}$ in the basis of eigenvectors of the Roe matrix $A_{j,k}$, i.e.
$U_{k}-U_{j}=\sum_{l=1}^{N}\alpha_{j,k;l}\vec{r}_{j,k;l},$
are computed by solving a linear system
$R_{j,k}\vec{\alpha}_{j,k}=U_{k}-U_{j}.$ (34)
Then one has
$D^{\pm}_{j,k}=\sum_{l=1}^{N}\alpha_{j,k;l}\lambda^{\pm}_{j,k;l}\vec{r}_{j,k;l},$
$\lambda^{+}=\frac{\lambda+\left|\lambda\right|}{2},\quad\lambda^{-}=\frac{\lambda-\left|\lambda\right|}{2}.$
On the other hand, the method based on the LF splitting may be oscillatory if the reconstructions are not performed in characteristic fields. To avoid this, the reconstructions are computed in practice as follows:
@@
Observe that, if the LF splitting is chosen and reconstructions in characteristic variables are performed, the right and left eigenvectors of the Roe matrices $A_{i,i+1}$ have to be computed. On the other hand, if the Upwind reconstruction is selected, the eigenvectors and eigenvalues of all the Roe matrices $A_{k,j}$ are required. Moreover, the linear system (34) has to be solved. Therefore, it is more computationally expensive for homogeneous problems. Nevertheless, it will be seen in Section 3 that the numerical treatment of the source term can compensate for this disadvantage.
Remark 1.
@@
$\frac{dU_{i}}{dt}=-\frac{1}{\Delta x}\left(F_{i+1/2}-F_{i-1/2}+\widehat{B}^{-}_{i+1/2}+\widehat{B}^{+}_{i-1/2}\right),$ (37)
where $F_{i+1/2}$ and $\widehat{B}_{i+1/2}$ are, respectively, standard WENO reconstructions of the flux function and the nonconservative terms, in which the nonlinear coefficients are the same.